rexco's before tax cost of debt is 6%, their cost of equity is 10%, their books show $10 million of debt and $30 million of equity. if their debt trades at par, their $1 million outstanding shares trade at $70/share, and their tax rate is 20%. what is their weighted average cost of capital
Answer #1

The weighted average cost of capital (WACC) is the cost of raising additional capital, with the weights representing the proportion of each source of financing that is used.

WACC = wdrd (1 - t) + were                                  

Where

wd        is the proportion of debt that the company uses when it raises new funds

rd         is the before-tax cost of debt = 6%

t           is the company’s tax rate = 20%

we        is the proportion of equity that the company uses when it raises new funds

re         is the cost of equity = 10%

We know that Market value of Debt = $ 10 million (as debt trades at par)

Market value of equity = number of outstanding shares * selling price of a share

= 1 million * $ 70 = $70 million

Therefore Weight of debt wd = $ 10 million/ (10+ 70) million = 10/80 = 0.125

Weight of equity = $70 million/ (10+ 70) million = 70/80 = 0.875

Therefore,

WACC = 0.125 * 6% *(1 – 0.20) + 0.875 * 10%

WACC = 0.6 + 8.75 = 9.35%